A spring is fixed to a wooden board with 2 masses suspend from it. The spring has stretched 6 cm. If m1 = 13 kg and m2 = 3 kg, what is the spring constant of the spring?


1
Expert's answer
2021-03-02T09:30:53-0500

Stretch in spring "x=6cm=0.06m"

Tw masses "m_1=13kg,m_2=3kg"


Let k be the spring constant -


Then Restoration force in the spring "F=kx=0.06k~~~~~~-(1)"


Then Weight of masses = "(m_1+m_2)g=(13+3)(9.8)"

"=16\times 9.8=156.8N~~~~~-(2)"


From eqn.(1) and (2) we have-

"0.06k=156.8\Rightarrow k=\dfrac{156.8}{0.06}=2613.33N/m"


hence The spring Constant is 2613.33N/m.